OpenStudy (melstutes):

What is the equation of the line parallel to the line 5x-2x = 7 through the ordered pair (6, 0)?

OpenStudy (anonymous):

this equation doesn't make any since

OpenStudy (melstutes):

5x- 2y = 7 sorry

OpenStudy (melstutes):

y = 5/2 x +7

OpenStudy (melstutes):

Is this right

OpenStudy (anonymous):

i dont think so because it does not go through (6,0)

OpenStudy (anonymous):

Ok @melstutes, you got your equation in the form y=mx+c, where c is y-intercept, x and y represent coordinates of a point and m is the gradient of the function. Now, if a line is parallel to that, it means it has the same gradient as the first function. So we already know that its m is 5/2. Next, we can substitute the coordinates of the point given to find the y-intercept. That is\[y=mx+c \rightarrow0=5/2(6)+c \rightarrow-15=c\] Therefore, its equation is\[y=5/2x-15\]